### v1

# Mingyu Yang, Mehdi Rezagholizadeh, Guihong Li , Vikram Appia, Emad Barsoum Advanced Micro Devices, Inc. (AMD)

Abstract
With the growing demand for deploying large language models (LLMs) across diverse applications, improving their inference efficiency is crucial for sustainable and democratized access. However, retraining LLMs to meet new user-specific requirements is prohibitively expensive and environmentally unsustainable. In this work, we propose a practical and scalable alternative: composing efficient hybrid language models from existing pre-trained models. Our approach, Zebra-Llama, introduces a family of 1B, 3B, and 8B hybrid models by combining State Space Models (SSMs) and Multi-head Latent Attention (MLA) layers, using a refined initialization and post-training pipeline to efficiently transfer knowledge from pre-trained Transformers. Zebra-Llama achieves Transformer-level accuracy with near-SSM efficiency using only 7–11B training tokens (compared to trillions of tokens required for pre-training) and an 8B teacher. Moreover, Zebra-Llama dramatically reduces KV cache size—down to 3.9%, 2%, and 2.73% of the original for the 1B, 3B, and 8B variants, respectively—while preserving 100%, 100%, and $>$97% of average zero-shot performance on LM Harness tasks. Compared to models like MambaInLLaMA, X-EcoMLA, Minitron, and Llamba, Zebra-Llama consistently delivers competitive or superior accuracy while using significantly fewer tokens, smaller teachers, and vastly reduced KV cache memory. Notably, Zebra-Llama-8B surpasses Minitron-8B in few-shot accuracy by 7% while using 8×8$\times$8 × fewer training tokens, over 12×12$\times$12 × smaller KV cache, and a smaller teacher (8B vs. 15B). It also achieves 2.6×2.6$\times$2.6 ×3.8×3.8$\times$3.8 × higher throughput (tokens/s) than MambaInLlama up to a 32k context length. We will release code and model checkpoints upon acceptance.
1 Introduction
Figure 1: Comparing 8B-scale models on average LM Harness score vs. KV cache size. Zebra-Llama (green) matches or exceeds baselines with smaller KV cache and fewer training tokens. Circle and square sizes indicate training tokens (billions for post-training, trillions for pre-training).
The exponential growth of deep learning applications has created an urgent demand for models that strike a balance between accuracy and computational efficiency—particularly in scenarios constrained by memory or limited hardware capabilities. Transformer-based models, despite their impressive performance across a range of tasks, are fundamentally limited by the quadratic complexity of their self-attention mechanisms and the substantial memory required to store key–value (KV) caches. These bottlenecks hinder their deployment in real-world applications, especially on edge devices or in latency-sensitive settings. At the same time, the rise of large language models (LLMs) has amplified the need for customization—that is, the ability to adapt pre-trained models to meet diverse user needs, hardware configurations, and application requirements. However, developing new LLMs from scratch for each target environment is prohibitively expensive and environmentally unsustainable. Traditional solutions such as model compression, neural architecture search (NAS), and pre-training new architectures offer potential pathways but suffer from significant limitations. Model compression often degrades quality while NAS or pre-training new models have substantial computational costs.
To overcome these challenges, a promising new paradigm has emerged: hybrid models that aim to reduce the computational cost of self-attention while maintaining generation quality. These architectures typically integrate efficient state-space models [1] or linear attention [2] with full attention mechanisms, leveraging the strengths of both. However, many recent hybrid approaches, including Samba [3] and Hymba [4], require extensive pre-training from scratch, which is computationally expensive. Others, such as MambaInLLaMA [5], experience notable performance degradation when attention is replaced too aggressively, often due to insufficient key–value cache compression or ineffective knowledge transfer from the base model.
Our goal in this work is to develop a more efficient and sustainable alternative: to compose highly efficient language models directly from existing pre-trained Transformers, avoiding the cost of full pre-training while retaining performance. Our approach, called Zebra-Llama, introduces a family of hybrid models (1B, 3B, and 8B) built on two complementary components: Multi-Latent Attention (MLA) [6], a low-rank attention mechanism that compresses memory usage without sacrificing quality under moderate compression; and Mamba2, a state-space model that eliminates KV caches entirely but performs poorly when used alone [5]. Specifically, we first initialize pure MLA and pure Mamba2 models from a pre-trained Transformer via a refined weight-mapping procedure. We then use Intermediate Layer Distillation (ILD) to align their internal representations with those of the original Transformer model, ensuring strong initialization. Finally, we strategically compose hybrid architectures from the refined MLA and Mamba2 variants using a sensitivity-aware strategy called SMART (Sensitivity Measure-Aware Replacement of Transformer layers) to select where each component is most effective. This process results in highly efficient models that retain Transformer-level quality with drastically reduced memory and compute requirements. Our Zebra-Llama family of 1B, 3B, and 8B hybrid models, achieving 25×25$\times$25 ×, 50×50$\times$50 ×, and 36×36$\times$36 × KV cache compression relative to their respective base Transformer models, while maintaining 100%, 100%, and >97% of the base model’s average zero-shot performance on the LM Harness evaluation benchmark. Our models also perform competitively on few-shot tasks. Notably, Zebra-Llama-8B improves the average few-shot accuracy over Minitron-8B by 7%, despite using 8×8$\times$8 × fewer training tokens, >12$\times$ smaller KV cache, and a smaller teacher model (8B vs. 15B). Additionally, our Zebra-Llama exhibits high inference speed. Compared to the existing work MambaInLlama, it achieves 2.6–3.8$\times$ higher throughput for sequences up to 32k tokens. This work introduces a practical route for building efficient, customizable LLMs from existing models. Our key contributions are:
  • Architecture: We propose a hybrid model combining MLA and Mamba2 layers, replacing classical Transformer blocks to reduce memory usage and address the quadratic bottleneck of attention.
  • Training: We develop an efficient post-training pipeline including refined initialization, Intermediate Layer Distillation, and SMART layer selection for hybrid model composition.
  • Empirical Results: Our models match or exceed Transformer-level performance with drastically reduced KV cache and significantly improved inference throughput.
2 Related Work
Hybrid Models
Prior work on hybrid models can broadly be categorized into two groups based on their training strategy: pre-training-based and post-training-based approaches. Pre-training-based methods, such as Jamba [7], Hymba [8], Samba [3] and Mamba-2-Hybrid [9], interleave heterogeneous hybrid layers during full-scale model training, allowing the development of hybrid models from scratch. While these models achieve strong performance, their training cost remains high, limiting accessibility and sustainability. In contrast, post-training approaches insert efficient modules into pre-trained Transformers, often leveraging knowledge distillation to transfer capabilities without full re-training, including MambaInLLaMA [5], MOHAWK [10], and Llamba [11]. Our work focuses on the post-training setting and pushes it further by targeting extremely efficiency—both in terms of training tokens and runtime memory. We introduce a systematic method to build hybrid models with minimal training costs, making them practical for broad deployment, especially where hardware resources are critically constrained."
Efficient Post-training
Several recent approaches specifically target post-training efficiency through careful initialization, distillation, and model compression. For example, MambaInLLaMA [5] replaces most self-attention blocks of a pre-trained Transformer with linear RNN layers. Through the initialization of RNN layer parameters from the pre-trained Transformer’s weights, followed by distillation-based fine-tuning, the hybrid model achieves comparable performance with the teacher model. MOHAWK [10] introduces a multi-stage cross-architecture distillation strategy (aligning attention “mixing” matrices, hidden states, and outputs) to transfer a Transformer’s knowledge into a Mamba-based model using a fraction of the original training data. Building on this, Llamba [11] scales the distilled recurrent architecture up to 8B parameters, attaining improved accuracy, while markedly improving inference throughput and memory usage for deployment on edge devices. Orthogonally, X-EcoMLA [12] “upcycles” a pre-trained Transformer’s attention into multi-head latent attention modules, jointly compressing key–value caches significantly with teacher-guided fine-tuning to preserve accuracy. Meanwhile, Minitron [13] compresses a 15B Nemotron Transformer by pruning its depth, width, and attention heads, then retrains on only about 2–3% of the original data via distillation; this yields 8B and 4B models that rival larger models’ performance without full re-training. Our method offers significant advantages over previous distillation techniques like Minitron, MambaInLlama, X-EcoMLA, and Llamba, which have each faced limitations in KV cache compression, training efficiency or maintaining the base models’ performance.
Figure 2: Overview of our hybrid model composition pipeline. The process consists of three stages: (1) Weight Initialization – we initialize pure Mamba2 and MLA models from a pre-trained Transformer via structured mapping; (2) Refined Initialization through Intermediate Layer Distillation (ILD) – we refine both models by aligning their internal representations with the base model on a small dataset; and (3) SMART Layer Selection – we compose the final hybrid model by selecting MLA and Mamba2 layers based on sensitivity analysis.
3 Methodology
Our methodology focuses on designing and training a hybrid model architecture that achieves strong performance with significantly enhanced efficiency. The overall approach consists of two key stages: (1) constructing an extremely efficient hybrid model, and (2) applying a lightweight yet effective training pipeline. An overview of the process is provided in Figure 2.
3.1 Composing Extremely Efficient Hybrid Model
To compose our hybrid model, we combine two complementary components: Mamba2 [14] and MLA [6] blocks. Each of these blocks contributes differently to efficiency and performance:
  • Mamba2 blocks are based on SSMs with zero KV cache usage, making them ideal for long-context or memory-constrained settings. However, they often underperform when used exclusively [9].
  • MLA blocks compress standard attention process to reduce KV cache requirements while maintaining high performance. Yet, excessive compression can lead to noticeable performance drops [12].
Our hybrid architecture interleaves Mamba2 and MLA layers to set a balance between minimal memory usage and strong predictive performance. The composition process consists of two key stages: (i) we first construct a pure Mamba2 model and a pure MLA model by replacing all attention blocks in the base Transformer with Mamba2 and MLA blocks, respectively, each initialized from the original pre-trained weights; (ii) we then apply intermediate layer distillation (ILD) to align the outputs of each layer in the Mamba2 and MLA models with those of the corresponding layers in the base model, thereby preserving the original model’s knowledge during the transition.
Following this refined initialization, we compose the final hybrid model using our sensitivity measure-aware replacement of transformer layers (SMART) layer selection strategy, which selects the optimal placement configurations of Mamba2 and MLA layers based on layer-wise sensitivity analysis.
3.1.1 Refined Initialization
To initialize the MLA and Mamba2 blocks in pure MLA and pure Mamba2 models derived from a pre-trained base model, we introduce an enhanced initialization strategy that goes beyond conventional weight transfer.
Background: Multi-Head Attention
Given an input hidden representation $H\in\mathbb{R}^{l\times d}$, MHA projects it into query, key, and value matrices using learned weights:
$Q=HW^{Q},\quad K=HW^{K},\quad V=HW^{V},$ (1)
where $W^{Q},W^{K},W^{V}\in\mathbb{R}^{d\times n_{h}d_{h}}$ and $n_{h}$, $d_{h}$ denote the number of heads and per-head dimension, $d$ is the hidden dimension, and $l$ represents the sequence length. The attention operation computes:
$A=\text{Softmax}\left(QK^{\top}/{\sqrt{d}}\right),\quad O=AVW^{O},$ (2)
with $W^{O}\in\mathbb{R}^{d\times d}$. During inference, MHA requires caching $K$ and $V$ for all past tokens, incurring a memory cost of $2n_{h}d_{h}l$.
Background: Multi-Head Latent Attention
MLA [6] introduces a low-rank compression scheme to reduce memory usage. Instead of storing $K$ and $V$ directly, they are compressed into a latent vector:
$C^{KV}=HW^{DKV},$ (3)
with $W^{DKV}\in\mathbb{R}^{d\times r_{kv}}$ and $r_{kv}\ll n_{h}d_{h}$. Keys and values are reconstructed via:
$K^{C}=C^{KV}W^{UK},\quad V^{C}=C^{KV}W^{UV},$ (4)
where $W^{UK}$ and $W^{UV}$ are up-projection matrices. MLA can also compress the query:
$C^{Q}=HW^{DQ},\quad Q^{C}=C^{Q}W^{UQ},$ (5)
where $C^{Q}\in\mathbb{R}^{l\times r_{q}}$ and $r_{q}\ll d$. To retain compatibility with RoPE, MLA decouples positional encoding using separate projections:
$Q^{R}=\text{RoPE}(C^{Q}W^{QR}),\quad K^{R}=\text{RoPE}(HW^{KR}),$ (6)
where $W^{QR}\in\mathbb{R}^{r_{q}\times n_{h}d_{r}}$ and $W^{KR}\in\mathbb{R}^{d\times d_{r}}$ The final queries and keys are then constructed as: $Q=[Q^{C};Q^{R}],\quad K=[K^{C};\text{repeat}(K^{R})]$, where $\text{repeat}(K^{R})$ denotes replication across heads. Overall, MLA reduces KV-cache size for inference from $O(n_{h}d_{h}l)$ to $O((r_{kv}+d_{r})l)$.
Initializing MLA from Pre-trained MHA
To construct an MLA-based hybrid model from a pre-trained Transformer, we upcycle its attention modules by reparameterizing them into low-rank latent attention. This conversion is initialized using a structured approach based on singular value decomposition (SVD), enabling knowledge transfer while minimizing performance loss [12]. The core MLA weights (i.e. query, key, and value projections) are initialized using low-rank approximations derived from pre-trained MHA parameters. For query weights, we apply SVD on $W^{Q}$:
$W^{Q}=U_{q}\Sigma_{q}V_{q}^{\top},$ (7)
and set $W^{DQ}=U_{q}$. The up-projection matrices $W^{UQ}$ and $W^{QR}$ are derived from $\Sigma_{q}V_{q}^{\top}$ by partitioning and reshaping into the appropriate query and RoPE dimensions. For key and value projections, we concatenate $W^{K}$ and $W^{V}$ and apply joint SVD:
$[W^{K},W^{V}]=U_{kv}\Sigma_{kv}V_{kv}^{\top}.$ (8)
We then set $W^{DKV}=U_{kv}$, and derive $W^{UK}$ and $W^{UV}$ by partitioning and reshaping $\Sigma_{kv}V_{kv}^{\top}$. The shared RoPE key projection $W^{KR}$ is initialized from the average $W^{K}$ across heads. We choose a constant rank $r_{q}$, $r_{kv}$ across all layers. Non-attention parameters such as the feedforward network and output projection $W^{O}$ are directly copied from the pre-trained model. Additional details of MLA initialization are provided in Appendix A.1.
Initializing Mamba2 from Attention Representations
It has been shown in [5] that linear attention can be reinterpreted as a special case of SSMs. This connection enables the initialization of Mamba2 blocks from pre-trained attention-based Transformers. In particular, the linearized form of attention without softmax resembles a linear RNN update:
$h_{t}=m_{t-1,t}h_{t-1}+K_{t}^{\top}V_{t},\quad y_{t}=\frac{1}{\sqrt{D}}Q_{t}h_{t},$ (9)
which parallels the linear RNN structure:
$h_{t}=A_{t}h_{t-1}+B_{t}x_{t},\quad y_{t}=C_{t}h_{t}.$ (10)
To bridge this connection, MambaInLLaMA [5] proposes to initialize the continuous-time Mamba2 SSM parameters ($A$, $B$, $C$) from the weights of the attention blocks. This includes discretizing the SSM over a learned sequence of sampling intervals $\Delta_{t}$ to match the temporal dynamics of attention. Specifically, $K_{t}^{\top}V_{t}$ in attention is mapped to $B_{t}x_{t}$ in Mamba2; $Q_{t}$ in attention plays the role of $C_{t}$; and the memory coefficient $m_{t-1,t}$ corresponds to $A_{t}$ in the recurrent update. Additional details of MLA initialization are provided in Appendix A.2.
Refined Initialization through Intermediate Layer Distillation (ILD)
After initializing the parameters of MLA and Mamba2 layers from the pre-trained Transformer, we further refine their weights through a lightweight ILD procedure on a small portion of the training data, akin to the second phase of MOHAWK [10]. This focused training aligns the internal representations between the MLA and Mamba2 layers and the pre-trained Transformer layers, ensuring a smoother start for downstream optimization and better preservation of the knowledge embedded in the original model. To perform ILD, we minimize the mean squared error (MSE) between the outputs of each pre-trained Transformer attention layer $h^{\text{Attn}}_{l}$ and the corresponding outputs of MLA ($h^{\text{MLA}}_{l}$) and Mamba2 ($h^{\text{M2}}_{l}$) layers. The ILD losses for Mamba2 and MLA are defined as:
$\begin{split}\mathcal{L}^{\text{\small{Mamba2}}}_{\text{ILD}}=\sum_{l\in\{1,2,...,L\}}\left\|h^{\text{Attn}}_{l}-h^{\text{M2}}_{l}\right\|_{2}^{2},~{}~{}\mathcal{L}^{\text{\small{MLA}}}_{\text{ILD}}=\sum_{l\in\{1,2,...,L\}}\left\|h^{\text{Attn}}_{l}-h^{\text{MLA}}_{l}\right\|_{2}^{2},\end{split}$ (11)
where $L$ denotes the total number of layers and $h_{l}$ refers to the output of the $l^{th}$ layer. Unlike MOHAWK, where the weights of MLP within each layer remain frozen during distillation, we allow training of all parameters in the Mamba2 and MLA layers. The refined initialization resulting from this ILD has proven to be crucial for enhancing the subsequent end-to-end distillation process, as evidenced in Section 4.3.3.
3.1.2 SMART: Sensitivity Measure-Aware Replacement of Transformer Layers
Figure 3: Layer sensitivity scores for Llama3.2-1B using 4096 samples from the validation dataset. Red markers indicate the MLA layer indices selected by our SMART strategy with $N=4$.
The final stage of our initialization process is to compose our hybrid model from the full MLA and full Mamba2 models. Let $N$ denote the desired total number of MLA layers. The set of layers assigned as MLA is represented by $\mathcal{L}_{\text{MLA}}\subseteq\{0,1,\ldots,L{-}1\}$. To minimize the performance gap between the hybrid and fully attention-based models, we introduce a Sensitivity Measure-Aware Replacement of Transformer Layers (SMART) strategy, which leverages empirical sensitivity analysis to guide the layer assignment. To measure the sensitivity of each layer, we begin with the full Mamba2 model after ILD. First, we compute the KL divergence between the full attention-based teacher model and the student model where all layers are Mamba2. Then, for each layer $i$, we construct a variant of the student model in which only the $i^{\text{th}}$ layer is replaced with MLA while the rest remain Mamba2, and measure the KL divergence against the teacher. The sensitivity gain $s_{i}$ for layer $i$ is defined as the reduction in KL divergence relative to the full Mamba2 model. The sensitivity score is formally computed as:
$\begin{split}s_{i}=\sum_{t=1}^{T}\mathbf{KL}\left[p(\cdot\mid y_{1:t},x,\theta_{T})\,\middle\|\,p(\cdot\mid y_{1:t},x,\theta)\right]-\mathbf{KL}\left[p(\cdot\mid y_{1:t},x,\theta_{T})\,\middle\|\,p(\cdot\mid y_{1:t},x,\theta^{\prime}_{i})\right],\end{split}$ (12)
where $\theta_{T}$ and $\theta$ denote the parameters of the teacher and full Mamba2 model, and $\theta^{\prime}_{i}$ corresponds to the variant with MLA inserted at layer $i$. In this equation, $y_{1:t}$ refers to model predictions up to time step $t$ and $x$ is the input sequence and $T$ is total number of decoding time steps. A higher score indicates that the $i^{\text{th}}$ layer plays a more critical role in aligning with the teacher and should thus be prioritized for MLA replacement due to its higher representational capacity. Figure 3 shows an example sensitivity profile for the Zebra-Llama 1B model after refined initialization, where we observe that the earliest and latest layers tend to exhibit higher sensitivity, while middle layers are typically less critical. While it may seem intuitive to simply select the top $N$ layers with the highest sensitivity scores, this strategy often leads to suboptimal layer placements—especially if it results in large gaps between MLA layers. To enforce a more structured distribution and preserve the most sensitive positions, we adopt the following three-step procedure:
  • Terminal Preservation: We divide the total $L$ layers into $N$ roughly uniform partitions. We take the first and last $\lfloor L/N\rfloor$ layers in the first and last partitions. From these two partitions, we select the layer with the highest sensitivity score as the first and last MLA layers, denoted $L_{1}^{\text{MLA}}$ and $L_{N}^{\text{MLA}}$, respectively. This ensures that the most sensitive edge layers are preserved.
  • Near-Uniform Intermediate Distribution: Given the range $[L_{1}^{\text{MLA}}+1,L_{N}^{\text{MLA}}-1]$, we aim to place the remaining $N-2$ MLA layers such that the gaps between consecutive MLA layers are as uniform as possible. We constrain the gap between adjacent MLA layers to lie within the range of min: $\lfloor\frac{L_{N}^{\text{MLA}}-L_{1}^{\text{MLA}}-N+1}{N-1}\rfloor$, and max: $\lceil\frac{L_{N}^{\text{MLA}}-L_{1}^{\text{MLA}}-N+1}{N-1}\rceil$. We enumerate all valid configurations $\{C_{j}\}_{j=1}^{M}$ that meet this spacing constraint, where each candidate $C_{j}$ defines a possible set of intermediate MLA layer indices.
  • Maximal Sensitivity Scores: For each valid configuration, we compute the total sensitivity score and choose the one with the highest cumulative score: $C^{*}=\underset{C_{j}}{\arg\max}\sum_{i\in C_{j}}s_{i}$.
Additional details and examples of this layer selection procedure are provided in Appendix B.
3.2 Efficient Training
After initialization and model composition, we follow an end-to-end knowledge distillation and DPO training stages to incrementally improve model accuracy and efficiency.
End-to-End Knowledge Distillation
This stage involves an end-to-end distillation with supervised fine-tuning (SFT):
$\mathcal{L}_{\theta}=\sum_{t=1}^{T}\mathbf{KL}[p(\cdot|y_{1:t},x,\theta_{T})||p(\cdot|y_{1:t},x,\theta)],$ (13)
where $\theta$ and $\theta_{T}$ are the parameters of the student model and the teacher model respectively. Such distillation step is crucial for transferring the rich, pre-trained knowledge from the teacher model.
Direct Preference Optimization (DPO)
In the final training stage, we perform DPO which is a binary cross entropy loss to adjust the preference of the student model. We set the distilled student model itself as the reference model as it makes the training much stabler.
4 Experiments and Results
4.1 Training Setup
All of our Zebra-Llama models are distilled from the Llama family: Llama3.2-1B-Instruct, Llama3.2-3B-Instruct, and Llama3.1-8B-Instruct for our experiments. For ILD and SFT, we use the same dataset as in [5] which includes multiple public datasets such as OpenHermes-2.5[15], GenQA[16], and Infinity-Instruct [17], with a total number of 6.8 billion tokens. The dataset is splited into $20\%$ and $80\%$ for ILD and SFT separately. We repeat the same training data more than one epoch to match the desired token budget. For DPO preference tuning, we adopt three datasets Llama3-ultrafeedback[18], orca_dpo_pairs[19], and ultrafeedback_binarized[20]. All models were trained on a single node equipped with eight AMD MI300 GPUs. Our training details are provided in Appendix A.4.
4.2 Performance Evaluation
Evaluation Tasks
We adopt the LM Harness Eval benchmark  [21] to perform zero-shot and few-shot evaluations on language understanding tasks, which includes ARC-Challenge (ARC) [22], ARC-Easy (ARE) [22],HellaSwag (HS) [23], MMLU (MM) [24], OpenBookQA (OB)  [25], PIQA [26], RACE (RA) [27], and WinoGrande (WG)  [28]. We provide more details for model evaluations in Appendix A.5.
Model and Setting Teacher Tokens Size KV Size ARC ARE HS MM OB PI RA WG Avg.
Llama3.2-1B-Inst - 9T 1.24B 100% 37.97 63.30 60.65 46.05 34.80 74.32 38.18 59.67 51.87
[…]
X-EcoMLA-1B ($r_{kv}=64$) 8B 7B 1.23B 9.37% 40.02 67.17 58.4 38.53 37.8 73.83 39.43 60.93 52.01
Llamba-1B 1B+70B 8B 1.41B 0% 37.12 65.36 61.31 38.11 36.8 73.78 37.61 60.62 51.34
[…]
Zebra-Llama-1B, 6MLA-10M2 8B 7B 1.28B 5.86% 43.94 67.51 60.46 38.21 41.2 73.23 37.61 61.17 52.92
Zebra-Llama-1B, 4MLA-12M2 8B 7B 1.28B 3.91% 42.15 67 58.86 37.96 40.4 73.18 37.89 58.88 52.04
Llama3.2-3B-Inst - 9T 3.21B 100% 46.08 67.93 70.38 60.34 36.4 75.79 40.86 67.25 58.13
[…]
X-EcoMLA-3B ($r_{kv}=816$) 3B 7B 3.21B 42.97% 48.38 70.37 72.41 57.51 38.2 76.28 46.41 68.11 59.71
X-EcoMLA-3B ($r_{kv}=128$)* 8B 7B 3.21B 9.37% 52.05 75.38 70.95 53.2 40.8 77.09 44.69 66.85 60.13
Llamba-3B 3B+70B 10B 3.66B 0% 45.65 73.78 73.31 52.32 42.4 78.02 40.1 70.01 59.45
[…]
Zebra-Llama-3B, 8MLA-20M2 8B 9B 3.36B 2.68% 51.96 75.97 72.38 48.16 42.8 77.64 43.54 65.67 59.77
Zebra-Llama-3B, 6MLA-22M2 8B 9B 3.39B 2.01% 49.74 74.62 71.11 44.21 43.8 77.37 42.11 65.43 58.55
Llama3.1-8B-Inst - 15T 8.03B 100% 54.86 79.55 79.23 68.13 43 80.9 44.69 73.88 65.53
[…]
Minitron-8B 15B 94B 8.3B 66.67% 52.73 79.5 77.4 62.95 45.2 81.39 39.71 72.69 63.95
X-EcoMLA-8B ($r_{kv}=128$)* 8B 7B 8.03B 9.37% 56.57 79.04 77.38 58.6 42.8 79.6 48.33 70.96 64.16
Llamba-8B 8B+70B 12B 8.32B 0% 53.84 79.71 76.25 60.29 44 79.16 40.38 72.77 63.3
[…]
Zebra-Llama-8B, 8MLA-24M2 8B 11B 8.38B 2.73% 57.17 79.46 77.43 55.4 44.6 79.71 44.31 71.43 63.69
Table 1: Zero-shot evaluation on the LM Harness Eval benchmark across eight tasks: ARC-Challenge (ARC), ARC-Easy (ARE), HellaSwag (HS), MMLU, OpenBookQA (OB), PI, RACE (RA), and WinoGrande (WG). All the teacher models are Llama3.1-8B/70B or Llama3.2-1B/3B execpt for Minitron (which is Nemotron-15B). The X-EcoMLA results are reproduced by ourselves.
Zero-shot Results
The results of our zero-shot evaluations are summarized in Table 1. We compare our Zebra-Llama with the base Llama models and other baselines based on distillation: MambaInLLaMA (Hybrid Mamba2-GQA)[5], X-EcoMLA (Pure MLA)[12], Llamba (Pure Mamba2) [11], and Minitron (Pruning) [13]. Besides the evaluation results, we list the teacher model size, number of training tokens, student model size (the number of parameters), and the KV cache size compared to the base Llama models. For the MLA layers in Zebra-Llama, we set $r_{kv}=128$ for the 1B and 3B models and $r_{kv}=160$ for the 8B models. For each model size, we tested various combinations of MLA and Mamba2 layers.
As shown in Table 1, compared to the base Llama models, our Zebra-Llama achieves extreme KV cache compression without noticeable performance drop. For 1B and 3B models, we achieves $3.91\%$ (25.6×25.6$\times$25.6 × compression) and $2.01\%$ (49.78×49.78$\times$49.78 × compression) KV cache size with even higher performance than the base Llama models. For the 8B models, we reach $5.47\%$ (18.3×18.3$\times$18.3 × compression) and $2.73\%$ (36.6×36.6$\times$36.6 × compression) KV cache size with only $0.87\%$ and $2.8\%$ performance drop. Note that for the 8B models, we achieve such compression ratio using only same size teachers. Further more, our Zebra-Llama achieves a better balance between KV cache compression and performance with much more efficient training (i.e., fewer training tokens or smaller teacher size) than other distillation-based methods. For example, compared with hybrid MambaInLlama, Zebra-Llama has similar performance with 12.79×12.79$\times$12.79 × and 24.88×24.$8$ smaller KV cache for 1B and 3B models. For 8B models, Zebra-Llama achieves 9.14×9.14$\times$9.14 × KV cache compression with only $1.6\%$ performance degradation by using a much smaller teacher (8B) and fewer training tokens (11B). Similarly, our method has roughly the same performance as the pure-MLA X-EcoMLA models, but with up to 3.4×3.4$\times$3.4 × less KV cache size. Moreover, compared to the pure Mamba model Llamba-8B, Zebra-Llama shows significantly better performance with both smaller teacher and fewer training tokens with minimal KV cache overhead. Moreover, we further compare Zebra-Llama with state-of-the-art hybrid models trained from scratch in Appendix C (see Table 8). While prior methods like SAMBA and Mamba-2-Hybrid rely on 1.5T–3.5T training tokens, our Zebra-Llama models achieve competitive or superior performance using only 7–11B tokens—representing a 214×–500× reduction in training data.
Few-shot Results
In Table 2, we report the results of few-shot evaluations on the Zebra-Llama-8B models under 25-shot ARC-Challenge (ARC), 10-shot HellaSwag (HS), 5-shot Winogrande (WG), 5-shot MMLU (MM), and 0-shot mc2 for Truthful-QA (TQ) tasks. Among all models, our Zebra-Llama achieves the best performance with only $5.47\%$ KV cache usage. The closest one to us is MambaInLlama-50% but it’s trained with 1.8×1.8$\times$1.8 × more tokens and has 9.14×9.14$\times$9.14 × more KV cache usage. For pure Mamba2 models such as MambaInLlama-8B-0% and Llamba-8B, they don’t use any KV cache but their performance is significantly worse than Zebra-Llama even with more training tokens and larger teachers.
Model and Setting KV % ARC HS MM WG TQ Avg.
Llama3.1-8B-Inst 100% 60.75 80.12 68.23 73.72 53.99 67.36
[…]
MambaInLlama-50% 50% 60.41 77.97 56.67 71.35 66.6 66.6
MambaInLlama-25% 25% 59.22 76.88 53.94 64.88 64.64 63.91
MambaInLlama-12.5% 12.5% 53.33 72.16 50.85 63.61 61.12 60.21
MambaInLlama-0% 0% 53.51 70.31 44.21 58.91 52.31 55.85
Llamba 0% 57.94 77.13 59.89 72.77 49.46 63.44
X-EcoMLA-8B ($128$)* 9.37% 59.64 76.9 58.73 71.43 60.86 65.51
[…]
Zebra-Llama, (8-24) 2.73% 60.41 76.11 55.06 71.11 61.01 64.74
Table 2: Few-shot evaluation on the LM Harness Eval benchmark across five tasks.
Throughput Evaluation
Figure 4: Inference throughput vs. output sequence length of various 8B-size models. We measure the throughput under batch size 1.
In Figure 4, we demonstrate the inference efficiency of Zebra-Llama by evaluating the throughput of different models employing distinct attention replacement strategies for Llama-3.1-8B. All experiments are conducted on a single AMD MI300X GPU. We fix the batch size and prompt length to 1 and measure the inference throughput across various generation lengths. As shown, our Zebra-Llama significantly outperforms the throughput of the MambaInLlama model (hybrid GQA-Mamba2). For long generation lengths like 32k, our method achieves throughput improvements of 2.6x and 3.79x with 8 MLA and 16 MLA layers compared to MambaInLlama with 8 GQA and 16 GQA layers. We note that MambaInLlama-8B-0% has slightly higher througput thanZebra-Llama; however, it has much worse accuracy scores accuracy, shown in Table 2.
4.3 Ablation Studies
In this section, we present a series of ablation studies aimed at justifying key design decisions in our approach. Specifically, we examine the impact of initialization strategies, the effectiveness of our SMART layer selection mechanism, the trade-offs between the number and size of MLA and Mamba2 layers, and the role of teacher model scaling.
4.3.1 Effect of Initialization Strategies
Figure 5: Performance of various initialization strategies after SFT for different model architectures. Generally, our proposed two-stage method achieves the highest average scores.
Figure 5 presents our assessment of various initialization methodologies through a comparison of three scenarios: Random weight initialization without ILD, Structured weight initialization without ILD, and the proposed Structured weight initialization with ILD. The results highlight that both structured weight initialization and ILD significantly boost SFT performance, especially when used together. For Mamba layers, ILD is crucial for aligning their outputs with the original models due to architectural differences from attention layers, providing the primary performance uplift. As for MLA layers, structured initialization can boost the accuracy significantly, which offers a strong foundation further refined by ILD.
4.3.2 Impact of SMART Layer Selection
Selection Strategy MLA Indices Total sen. Avg. Score
Uniform #1 [0,4,8,12] 1787.9 48.84
Uniform #2 [3,7,11,15] 1055.8 49.72
Max score [0,1,2,14] 2672.5 48.98
[…]
Possible middle #2 [0,5,9,14] 2297.5 49.95
SMART (ours) [0,5,10,14] 2500.2 50.15
Table 3: SFT results with different layer selection strategies on Llama-3.2-1B model.
In Table 3, we demonstrate the benefits of our SMART layer selection strategy by comparing with other layer selections for our Zebra-Llama-1B with 4 MLA layers, by using the configs shown in Figure 3 as an example. First, by comparing the first two layer selections sets and our optimal solutions, we can conclude that Terminal Preservation strategy (i.e., always selecting layers from the first and last few layers) are important for preserving model accuracy. Second, we also testify the naive greed selection greedy method (the third row); the results shows that uniformly distributing the MLA layers has a major contribution to final performance as well, which verified the effectiveness of our Near-Uniform Intermediate Distribution strategy. Third, by analyzing the last three selections, together with Terminal Preservation and Near-Uniform Intermediate Distribution , Maximal Sensitivity Scores is indeed a good indicator for final accuracy score. In short, these outcomes underscore how each of the three pillars–when appropriately combined–contributes critically to the SMART layer selection strategy’s effectiveness.
4.3.3 Tradeoff between number of MLA layers and $r_{kv}$
Figure 6: Tradeoff between KV rank $r_{kv}$ and the number of MLA layers when fixing the total KV-cache size of given model.
The KV-cache size in Zebra-Llama is determined by two factors: the number of MLA layers and the KV rank $r_{kv}$ of each MLA layer. Figure 6 presents our findings from varying these two factors for our 1B model while maintaining a constant total KV-cache size. We observed that for significant KV-cache compressions, such as 12.8×12.8$\times$12.8 × and 25.6×25.6$\times$25.6 ×, optimal performance typically occurs with a moderate number of MLA layers, around six, coupled with an intermediate $r_{kv}$. Deviating from this balance will hurt the model’s performance. On one hand, increasing the number of MLA layers excessively, like to 16 layers at a 25.6×25.6$\times$25.6 × compression, forces the $r_{kv}$ per layer to become very small (e.g., $r_{kv}$ =8), which significantly degrades performance. On the other hand, reducing the number of MLA layers too much also leads to a decline in performance, since the model will then consist almost entirely of Mamba2 layers, which typically have a higher capacity gap with MHA layers than MLA layers do.
4.3.4 Scaling the Teacher
Effective knowledge transfer from a well-chosen, pre-trained teacher model is crucial for our method’s success. As shown in Table 4, appropriately scaled teachers significantly enhance student model performance though with a diminishing return when gradually increasing size of teacher models. This is due to the "capacity gap" phenomenon in knowledge distillation, where a smaller student model struggles to fully absorb and generalize the complex teacher’s representations when the teacher model far exceeds student mimicry capabilities [29, 30]. Therefore, it’s vital to select a teacher model that is sufficiently powerful to provide rich knowledge but not so extremely larger than the student model, thereby balancing distillation efficacy with student-teacher compatibility. Investigating adaptive teacher scaling or multi-stage distillation offers future solutions to these capacity limitations.
Model and Setting Teacher Size ARC ARE HS MM OB PI RA WG Avg.
Llama3.2-1B-Inst - 37.97 63.30 60.65 46.05 34.80 74.32 38.18 59.67 51.87
[…]
Zebra-Llama-1B, 4MLA-12M2 3B 39.51 62.79 57.61 37.94 38.20 72.52 36.94 56.59 50.26
Zebra-Llama-1B, 4MLA-12M2 8B 42.15 67 58.86 37.96 40.4 73.18 37.89 58.88 52.04
Zebra-Llama-1B, 4MLA-12M2 70B 43.17 69.57 57.77 39.45 38.80 72.80 38.09 59.83 52.44
Llama3.2-3B-Inst - 46.08 67.93 70.38 60.34 36.4 75.79 40.86 67.25 58.13
[…]
Zebra-Llama-3B, 8MLA-20M2 8B 51.54 75.55 71.52 47.12 43.6 77.2 42.68 65.9 59.39
Zebra-Llama-3B, 8MLA-20M2 70B 51.96 77.23 69.46 48.32 43.4 76.01 43.35 65.19 59.37
Llama3.1-8B-Inst - 54.86 79.55 79.23 68.13 43 80.9 44.69 73.88 65.53
[…]
Zebra-Llama-8B, 8MLA-24M2 70B 58.53 80.72 76.64 53.82 45.4 80.03 43.06 69.61 63.48
Table 4: Impact of scaling up teacher size on model performance trained on 7B tokens. Except for teacher size, we use the same training configurations for the same size of student models.
5 Conclusion
In this work, we addressed the growing need for efficient LLMs by proposing a practical and scalable framework for composing hybrid models from existing pre-trained Transformers. Motivated by the cost and environmental impact of retraining large models for downstream use, we introduced Zebra-Llama, a family of 1B, 3B, and 8B hybrid models built using SSMs and MLA layers. We developed an effective initialization scheme and a post-training knowledge transfer pipeline that enabled these hybrid models to inherit capabilities from larger teacher models with minimal additional training. Our approach significantly reduced memory requirements while preserving or exceeding the accuracy of strong baselines, demonstrating the viability of post-training hybridization as an efficient alternative to full retraining.
Limitations and Future Work
Our work opens several directions for future research. One limitation is the current focus on a single family of base models; in future work, we plan to explore hybridization strategies across diverse LLM architectures, including multimodal models and Mixture-of-Experts (MoE) frameworks. We also aim to scale our training pipeline to support larger models and extend our method to reasoning-intensive architectures. Another limitation lies in our reliance on strong teacher models for knowledge transfer. In some cases, a sufficiently large teacher may not be available for a given base model. Even when such teachers exist, using a separate teacher model for distillation adds to the memory and resource requirements during training. Addressing these issues—such as through teacher-free distillation or efficient self-distillation—remains an important avenue for future work.
Appendix A More Experimental Details for Zebra-Llama
A.1 Structured MLA Initialization
# MHA weights: W_Q, W_K, W_V
# MLA weights: W_DQ, W_UQ, W_QR, W_DKV, W_UK, W_KR, W_UV
# Initialization of W_DQ, W_UQ, and W_QR
U_q, sigma_q, V_q = svd(W_Q)
W_DQ = U_q
W_UQR_bar = (sigma_q @ V_q).view(r_q, n_h, d_h)
W_UQ = W_UQR_bar[:, :, :d_qk].view(r_q, n_h*d_qk)
W_QR = W_UQR_bar[:, :, -d_r:].view(r_q, n_h*d_r)
# Initialization of W_DKV, W_UK, W_KR, W_UV
U_kv, sigma_kv, V_kv = svd(torch.cat((W_K, W_V), -1))
W_DKV = U_kv
W_K_avg = W_K.view(d, n_kv, d_h).mean(1)
W_KR = W_K_avg[:, -d_r:]
W_UKV = sigma_kv @ V_kv
W_UK_bar = W_UKV[:, :d_h*n_kv].view(r_kv, n_kv, d_h)
W_UK = W_UK_bar[:,:,:d_qk].view(r_kv, n_kv*d_qk)
W_UV = W_UKV[:, d_h*n_h:]
Algorithm 1 Python-like pseudocode of the proposed SVD initialization for MLA.
Our SVD-based MLA layer initialization follows the methodology outlined in X-EcoMLA [12] for Multi-Head Attention (MHA). However, for Generalized Question Answering (GQA) models like the Llama 3 series, our approach diverges slightly by keeping the number of key/value heads from the base model for MLA while X-EcoMLA forces the number of key/value heads to be the same as the number of query head. With such modification, we notice trivial performance difference but with slightly fewer number of parameters.
Essentially, the matrices from the base MHA/GQA module could be expressed as $W^{Q}\in\mathbb{R}^{d\times n_{h}d_{h}}$, and $W^{K},W^{V}\in\mathbb{R}^{d\times n_{kv}d_{h}}$, where $d$ denotes the hidden dimension, $n_{h}$ denotes the number of attention heads, $n_{kv}$ denotes the number of key/value heads. For MHA, we have $n_{h}=n_{kv}$ while for GQA we have $n_{h}>n_{kv}$.
All the matrices we need to initialize in MLA could be expressed as $W^{DKV}\in\mathbb{R}^{d\times r_{kv}}$, $W^{UK}\in\mathbb{R}^{r_{kv}\times n_{kv}d_{qk}}$, $W^{UV}\in\mathbb{R}^{r_{kv}\times n_{kv}d_{v}}$, $W^{DQ}\in\mathbb{R}^{d\times r_{q}}$, $W^{UQ}\in\mathbb{R}^{r_{q}\times n_{h}d_{qk}}$, $W^{KR}\in\mathbb{R}^{d\times d_{r}}$, and $W^{QR}\in\mathbb{R}^{r_{q}\times n_{h}d_{r}}$, where $r_{kv}$ represents the latent dimension for key/value, $d_{qk}$ denotes the head dimension for query/key, $d_{v}$ denotes the head dimension for value, $r_{q}$ denotes the latent dimension for query, and $d_{r}$ represents the dimension for RoPE embeddings. For all the experiments, we keep $d_{v}=d_{h}$ and $d_{qk}+d_{r}=d_{h}$.
Initialization of $W^{DQ}$, $W^{UQ}$, and $W^{QR}$
Given that query compression in MLA can be viewed as a low-rank approximation of attention layers, we initially perform SVD on the pre-trained weight matrix $W^{Q}$:
$W^{Q}=U_{q}\Sigma_{q}V^{T}_{q},$ (14)
where $U_{q}\in\mathbb{R}^{d\times r_{q}}$, $\Sigma_{q}\in\mathbb{R}^{r_{q}\times r_{q}}$, and $V_{q}\in\mathbb{R}^{d_{h}n_{h}\times r_{q}}$. For constructing the up-projection matrices, we reshape the product $\Sigma_{q}V^{T}_{q}$ to form $\overline{W}^{UQR}\in\mathbb{R}^{r_{q}\times n_{h}\times d_{h}}$ and subsequently split this tensor at the last dimension into $W^{UQ}$, containing the first $d_{qk}$ elements, and $W^{QR}$, containing the remaining $d_{r}$ elements. The down-projection matrix $W^{DQ}$ is directly assigned with $U_{q}$. This method of initial assignment is formulated as:
$W^{DQ}=U_{q},\;\;W^{UQ}=\text{reshape}(\overline{W}^{UQR}[:,:,:d_{qk}]),\;\;W^{QR}=\text{reshape}(\overline{W}^{UQR}[:,:,-d_{r}:]),$ (15)
[…]
Initialization of $W^{DKV}$, $W^{UK}$, $W^{UV}$, and $W^{KR}$
The initialization of the MLA weights associated with keys and values is more complicated because of the decoupled RoPE mechanism. First, a joint SVD is performed on the concatenated $W^{K}$ and $W^{V}$:
$[W^{K},W^{V}]=U_{kv}\Sigma_{kv}V^{T}_{kv},$ (16)
where $U_{kv}\in\mathbb{R}^{d\times r_{kv}}$, $\Sigma_{kv}\in\mathbb{R}^{r_{kv}\times r_{kv}}$, and $V_{kv}\in\mathbb{R}^{2d_{h}n_{kv}\times r_{kv}}$. For the down-projection matrix $W^{DKV}$, we directly set it to $U_{kv}$. To derive the up-projection matrix $W^{UV}$ and $W^{UK}$, we first set $W^{UKV}=\Sigma_{kv}V^{T}_{kv}$. Since we have $d_{v}=d_{h}$, we simply extract the last $d_{h}n_{kv}$ columns of $W^{UKV}$ as $W^{UK}$. For $W^{UK}$, we first extract the first $d_{h}n_{kv}$ columns of $W^{UKV}$ and reshape them into $\overline{W}^{UK}\in\mathbb{R}^{r_{kv}\times n_{kv}\times d_{h}}$. Then, we select the first $d_{qk}$ elements along the last dimension of $\overline{W}^{UK}$ and reshape it back to obtain $W^{UK}$. In general, this process can be expressed as:
$W^{DKV}=U_{kv},\;\;W^{UV}=W^{UKV}[:,:n_{kv}d_{h}],\;\;W^{UK}=\text{reshape}(\overline{W}^{UK}[:,:,:d_{qk}]).$ (17)
In the final step, the initialization of the RoPE key embedding matrix $W^{KR}$ requires a distinct approach, given that all attention heads in MLA utilize the identical RoPE key embedding. First, the average key projection matrix $W^{K}_{avg}\in\mathbb{R}^{d\times d_{h}}$ is calculated for all attention heads. Then, the final $d_{r}$ columns are extracted for initializing $W^{KR}$, formulated as follows:
$W^{KR}=W^{K}_{avg}[:,-d_{r}:].$ (18)
The detailed initialization algorithm can be found in Algorithm 1.
A.2 Structured Mamba2 Initialization
The structured initialization process of Mamba2 layers follows precisely to the solution of MambaInLlama [5]. As outlined in Section 3.1.1, excluding the softmax operation in attention shows a direct one-to-one mapping between $B_{t}$, $x_{t}$, and $C_{t}$ in linear RNN and $K_{t}$, $V_{t}$, and $Q_{t}$ in attention layers. In the Mamba2 framework, $B_{t}$, $x_{t}$, and $C_{t}$ for the continuous-time SSM are derived from input $o_{t}$ via passing through a MLP followed by a 1d convolution layer. The MLP is replaced directly with pre-trained transformer layer weights as follows:
$x_{t}=W^{V}o_{t}\in\mathbb{R}^{b\times n_{kv}h_{d}},B_{t}=W^{K}o_{t}\in\mathbb{R}^{b\times n_{kv}h_{d}},C_{t}=W^{Q}o_{t}\in\mathbb{R}^{b\times n_{h}h_{d}},$ (19)
where $b$ signifies the batch size. Subsequent to this, $x$, $B$, and $C$ are processed through the 1d convolutional layer for temporal fusion before undergoing discretization in Mamba SSM. It is important to highlight that for GQA and MQA scenarios where $n_{kv}<n_{q}$, $x_{t}$ and $B_{t}$ are replicated after the convolution to ensure that $x_{t}$, $B_{t}$, and $C_{t}$ share identical dimensions. Other parameters, such as $A$ and $\Delta_{t}$, adhere to the original initialization procedure within the Mamba2 layers.
A.3 Model Structure Details
Detailed architectures of our 1B, 3B, and 8B models, as presented in Table 1, are comprehensively outlined in Table 5. This includes MLA layer selections, parameters specific to the MLA layers, and the overall model size. Layers not categorized as MLA are Mamba2 layers, which follows the configuration used in MambaInLlama [5].
Method MLA Indices $r_{kv}$ $r_{q}$ $d_{qk}$ $d_{r}$ Model Size
Zebra-Llama-1B, 8MLA8M2 [0,2,4,8,10,12,14] 128 1344 32 32 1.27B
Zebra-Llama-1B, 6MLA10M2 [0,2,5,8,11,14] 128 1344 32 32 1.28B
Zebra-Llama-1B, 4MLA12M2 [0,5,10,14] 128 1344 32 32 1.28B
[…]
Zebra-Llama-3B, 8MLA20M2 [0,4,8,12,16,20,24,27] 128 1536 64 64 3.36B
Zebra-Llama-3B, 6MLA22M2 [0,5,11,17,22,27] 128 1536 64 64 3.39B
[…]
Zebra-Llama-8B, 8MLA24M2 [0,4,8,13,18,23,27,31] 160 2048 64 64 8.38B
Table 5: Configurations of Zebra-Llama models’ architecture.
A.4 Training Details
In Table 6, we present the training configurations for our Zebra-Llama series models, including the number of tokens, batch size, learning rate, and total training time. All experiments are conducted on a single node equipped with eight AMD MI300 GPUs, each featuring 192GB of memory. We apply a learning rate warmup over the first $1\%$ of training data, followed by cosine annealing. The models are optimized using AdamW, with hyperparameters set to $\beta=(0.9,0.8)$. Additionally, all models process input sequences of length 2048 through sample packing.
Stage Model Teacher Tokens BS LR Time (H)
ILD Zebra-Llama-1B-MLA Llama3.2-1B 1.36B 96 8e-5 1.8
ILD Zebra-Llama-1B-Mamba2 Llama3.2-1B 1.36B 96 8e-5 1.7
ILD Zebra-Llama-3B-MLA Llama3.2-3B 1.36B 96 8e-5 3.9
ILD Zebra-Llama-3B-Mamba2 Llama3.2-3B 1.36B 96 8e-5 4.4
ILD Zebra-Llama-8B-MLA Llama3.1-8B 1.36B 48 4e-5 9.2
ILD Zebra-Llama-8B-Mamba2 Llama3.1-8B 1.36B 48 4e-5 10.3
[…]
SFT Zebra-Llama-3B Llama3.1-8B 7.44B 96 4e-5 $\approx 31.2$
SFT Zebra-Llama-8B Llama3.1-8B 9.44B 64 4e-5 $\approx 78.1$
[…]
DPO Zebra-Llama-3B N/A 0.2B 32 5e-7 $\approx 1.2$
DPO Zebra-Llama-8B N/A 0.2B 32 5e-7 $\approx 2.3$
Table 6: Training hyper-parameters of Zebra-Llama models.
A.5 Evaluation Details
We evaluate all models using the lm-evaluation-harness library (commit from the big-refactor branch) following the task-specific few-shot configurations defined by the Open LLM Leaderboard. For zero-shot evaluation, we report performance across a broad suite of language understanding tasks: MMLU, HellaSwag, PIQA, ARC-Easy, ARC-Challenge, Winogrande, OpenBookQA, and RACE. Evaluations are performed using the command-line interface with ROCm-enabled devices and a batch size of 16. For few-shot runs targeting leaderboard comparisons, we use the officially recommended number of shots per task (e.g., 25-shot for ARC-Challenge, 10-shot for HellaSwag, 5-shot for Winogrande and MMLU, 0-shot mc2 for Truthful-QA(TQ))  [31]. We report average scores across tasks, following the same protocol as prior work.
Appendix B SMART Layer Selection Algorithms
Algorithm 2 Pseudo code: SMART—Structured MLA Layer Selection via Sensitivity Scores
1:Sensitivity scores $\{s_{1},s_{2},\ldots,s_{L}\}$, number of MLA layers $N$
2:Selected MLA layer indices $\{L_{1}^{\text{MLA}},\ldots,L_{N}^{\text{MLA}}\}$
3:// Terminal Preservation
4:Divide $L$ layers into $N$ equal partitions
5:$L_{1}^{\text{MLA}}\leftarrow$ highest-sensitivity layer in first partition of layers $\left\{i,i\in\left[1,\left\lfloor\frac{L}{N}\right\rfloor\right]\right\}$
6:$L_{N}^{\text{MLA}}\leftarrow$ highest-sensitivity layer in last partition of layers $\left\{i,i\in\left[L-\left\lfloor\frac{L}{N}\right\rfloor+1,L\right]\right\}$
7:// Near-Uniform Intermediate Distribution
8:// Define Valid Intermediate Layer Range
9:Let $r_{\text{start}}\leftarrow L_{1}^{\text{MLA}}+1$
10:Let $r_{\text{end}}\leftarrow L_{N}^{\text{MLA}}-1$
11:Let $R\leftarrow$ list of candidate intermediate layers in $[r_{\text{start}},r_{\text{end}}]$
12:// Compute Allowable Gap Bounds
13:Let $T\leftarrow L_{N}^{\text{MLA}}-L_{1}^{\text{MLA}}-N+1$
14:Let $g_{\min}\leftarrow\left\lfloor\frac{T}{N-1}\right\rfloor$
15:Let $g_{\max}\leftarrow\left\lceil\frac{T}{N-1}\right\rceil$
16:// Enumerate All Valid Configurations
17:Initialize empty list of configurations $\mathcal{C}\leftarrow[]$
18:for all combinations of $N-2$ layers from $R$ do
19:     Sort the selected layers in ascending order to form $C_{j}$
20:     Let $G_{j}\leftarrow$ list of gaps between consecutive layers in $\{L_{1}^{\text{MLA}}\}\cup C_{j}\cup\{L_{N}^{\text{MLA}}\}$
21:     if all gaps in $G_{j}$ satisfy $g_{\min}\leq\text{gap}\leq g_{\max}$ then
22:         Append $C_{j}$ to $\mathcal{C}$
23:     end if
24:end for
25:// Maximal Sensitivity Scores
26:$C^{*}\leftarrow\arg\max_{C_{j}}\sum_{i\in C_{j}}s_{i}$
27:return $\{L_{1}^{\text{MLA}}\}\cup C^{*}\cup\{L_{N}^{\text{MLA}}\}$
We provide the pseudo code for SMART in Algorithm 2. Besides, we provide three examples for the layer selection process of our Zebra-Llama 1B models following SMART. The examples are based on the sensitivity analysis shown in Figure 3, whose actually values are listed in Table 7.
Layer Index Sensitivity Score Layer Index Sensitivity Score
0 1185.06 8 238.1
1 382.73 9 120.56
2 480.68 10 323.23
3 350.95 11 228.9
4 196.03 12 168.69
5 367.82 13 233.87
6 250.45 14 624.03
7 114.44 15 361.47
Table 7: Actually sensitivity score in Figure 3.
Example 1: Zebra-Llama-1B with $N=4$
Terminal Preservation: $L^{MLA}_{1}=0$, $L^{MLA}_{N}=14$ Near-Uniform Intermediate Distribution: Define intermediate layer range: $r_{start}=1$, $r_{end}=13$, $R=\{1,2,\ldots,13\}$. Compute allowable gap bounds: $T=11$, $g_{min}=3$, $g_{max}=4$. Enumerate valid Configurations: $\mathcal{C}=\{\{4,9\},\{5,9\},\{5,10\}\}$. Maximal Sensitivity Scores: Calculate total sensitivity score: * $s_{4}+s_{9}=316.59$, * $s_{5}+s_{9}=488.38$, * $s_{5}+s_{10}=691.05$ Layers with maximal score: $C^{*}=\{5,10\}$. Return $\{0,5,10,14\}$.
Example 2: Zebra-Llama-1B with $N=6$
[…]
Example 3: Zebra-Llama-1B with $N=8$
Terminal Preservation: $L^{MLA}_{1}=0$, $L^{MLA}_{N}=14$ Near-Uniform Intermediate Distribution: Define intermediate layer range: $r_{start}=1$, $r_{end}=13$, $R=\{1,2,\ldots,13\}$. Compute allowable gap bounds: $T=7$, $g_{min}=1$, $g_{max}=1$. Enumerate Valid Configurations: $\mathcal{C}=\{\{2,4,6,8,10,12\}\}$. Only one valid configuration. Return $\{0,2,4,6,8,10,12,14\}$.
Appendix C Comparison with Pre-training Methods
In Table 8, we include more comparisons of our Zebra-Llama with other hybrid models based on pre-training instead of distillation. It can be observed that our method could achieve competitive performance as the pre-trained baselines 214×500×214$\times$-500$\times$214 × - 500 × fewer training tokens, demonstrating our advantage in the training efficiency. For instance, Zebra-Llama (16MLA–16M2) matches or exceeds the accuracy of Mamba-2-Hybrid (8.66B, 3.5T tokens) and SAMBA (1.7B, 3.5T tokens), while using only 11B tokens and smaller model sizes. This highlights the efficiency and practicality of our post-training hybrid composition strategy, as it achieves high performance with a fraction of the training budget required by scratch-trained models.
Model and Setting Tokens Size ARC ARE HS MM OB PI RA WG
Mamba-2-Hybrid 3.5T 8.66B - - - -/47.7 77.23/- - - - - - - -/77.68 51.46 - - - -/42.8 79.65/- - - - 39.71 71.27
SAMBA 3.5T 1.7B 48.21/- - - - 79.25/- - - - 49.74/- - - - 48.01 37.20/- - - - 77.10/- - - - - 72.93
SAMBA 3.5T 1.3B - 58.21/- - - - - - - -/54.73 - - 72.36/- - - - - 55.72
Hymba 1.5T 1.5B 42.32/- - - - 74.54/- - - - 53.55/- - - - - - 76.66/- - - - - 66.61
[…]
Zebra-Llama, 14MLA-14M2 9B 3.27B 50/52.65 80.47/76.35 53.52/72.43 51.97 31.8/44.4 76.61/76.99 46.99 67.8
Zebra-Llama, 16MLA-16M2 11B 8.19B 57.17/58.96 83.59/79.92 57.82/77.73 57.18 35.20/44.6 79.65/80.2 48.71 72.38
Table 8: Comparing our Zebra-Llama with state-of-the-art hybrid models that are trained from scratch. We report two accuracy scores for each model, i.e, accuracy and normalized accuracy (acc / acc_norm). The missing results (‘-’) in the table means the results are not reported in original paper.
Appendix D More Inference Evaluations
Figure 7: Inference throughput vs. output sequence length of various 3B-size models. We measure the throughput under batch size 1.
In this section, we provide more inference evaluations on our Zebra-Llama 3B and 8B models. In Figure 7, we include the throughput evaluations for 3B models following the same setup with Figure 4 where the prompt length and batch size are fixed to 1. Similarly, our Zebra-Llama exhibits much better inference throughput than the hybrid GQA-Mamba model MambaInLlama. With 8k generation length, our Zebra-Llama is 2.26×2.26$\times$2.26 × faster than MambaInLlama with 14 MLA/GQA layers and 6.33×6.$3$ faster than MambaInLlama with 6 MLA/GQA layers.
In Figure 8 and 9, we include the inference evaluations for 8B models iunder a setting where the prompt length is fixed at 1, the output length at 8192, and the batch size gradually increases from 8 to 256. All measurements are recorded on a single AMD MI300 GPU with 192GB of memory, using the generation function from the mamba_ssm library. Our Zebra-Llama models demonstrate significant throughput gains compared to Llama, pure MLA models, and hybrid GQA-Mamba models. Although pure Mamba models achieve faster inference speeds than Zebra-Llama, they tend to exhibit poorer overall performance, as shown in Tables 1 and 2. Additionally, the KV cache compression in our Zebra-Llama models translates to notable memory savings, as illustrated in Figure 9. Models incorporating GQA layers experience a sharp increase in memory usage as the batch size grows. In contrast, MLA-based approaches, such as X-EcoMLA and our Zebra-Llama models, demonstrate superior memory efficiency, exhibiting only a marginal increase in memory consumption compared to the pure Mamba baseline.
Figure 8: Inference throughput vs. batch size of various 8B-size models. We measure the throughput with output sequence length 8192.
Figure 9: Peak memory usage during inference vs. batch size of various 8B-size models. We measure the memory footprint with output sequence length 8192.